OpenStudy (anonymous):

What is the product of the two real solutions of the equation 2x = 3 - x2 A) 3 B) 2 C) -2 D) 6 E) -3

OpenStudy (anonymous):

6 min left 0.0

OpenStudy (juanpablojr):

2*-3=3-(-3)*2 so that would be E see what happens

OpenStudy (anonymous):

uh @ganeshie8 is it E????

OpenStudy (here_to_help15):

r u done after this?????

OpenStudy (anonymous):

I need 1 more and done

OpenStudy (here_to_help15):

post the next 1 quick

OpenStudy (here_to_help15):

how many mins

OpenStudy (anonymous):

We can generalize our result to quadratic equations of the form \(p(x) = 0\) where \(p(x) = ax^2+bx+c = a(x^2+\frac bax+\frac ca)\) \(p(x) = a(x-x_1)(x-x_2) = a(x^2-x(x_1+x_2)+x_1x_2))\) We see that \(x_1x_2 = \cfrac ca\) for any quadratic equation! In our case we can write \(x^2+2x-3 = 0\), so \(a = 1,\;c = -3\) and we have \(x_1x_2 = \frac{-3}{1} = -3\), as already said.